# Rubber Friction: Theory, Mechanisms, and Challenges

### v2

@@ 5.1 Analytical theory of rubber friction on rough surfaces
For the relatively stiff rubber compounds used in practical applications such as tires and seals, there is consistent experimental evidence that crack-opening mechanisms contribute little to the overall sliding or rolling friction. This is primarily due to the multiscale roughness of the surface, which strongly suppresses adhesion [58]. In Sec. 8 we will show that for rubber compounds which exhibiting strong adhesion on smooth surfaces under dry conditions, but no adhesion in water, display nearly identical sliding friction on rough surfaces under both dry and wet conditions [26]. Similar observations have been made by Greenwood and Tabor [18].
@@ 5.2 Comparing the theory with experimental results
The results in Fig. 20 demonstrate that, in most cases, at low sliding speeds water is almost completely expelled from the asperity contact regions, resulting in a friction coefficient similar to that observed under dry conditions. However, at much higher sliding speeds, there is insufficient time for the water film to be fully squeezed out, leading to a significant reduction in friction. At very high speeds, hydroplaning may occur, and the friction becomes negligible.
@@
One important quantity that characterizes rubber is its glass transition temperature, $T_{\rm g}$, defined as the temperature at which ${\rm tan}\,\delta$ reaches its maximum for a given reference frequency (we use $\omega=0.01\ {\rm s}^{-1}$). Due to the WLF temperature–frequency relationship, $T_{\rm g}$ also provides information about the frequency at which ${\rm tan}\,\delta$ is maximal.
@@ 5.3 Dynamical friction
Figure 27: Slip at the onset of sliding when the driving force is applied on one side of the rubber block. The region where slip occurs gradually expands as the system approaches the onset of global slip, at which point the entire contact area is sliding. The frictional shear stress in regions where the slip displacement $u>\lambda_{\rm asp}$ is equal to the kinetic frictional shear stress $\tau_{\rm k}$. Consequently, the breakloose friction force is smaller than $F_{\rm s}=A\tau_{\rm s}$ and approaches $F_{\rm k}=A\tau_{\rm k}$ as the elastic modulus decreases.
@@
Consider a “kinetic” frictional shear stress $\tau_{\rm k}$ acting on the contact area $A$ between a block and substrate during sliding. After a stationary contact duration $t_{\rm s}$, the shear stress at the new onset of sliding becomes $\tau_{\rm s}$, which we assume to be larger than $\tau_{\rm k}$, e.g., due to thermally activated interfacial bond formation. To reduce the shear stress to its kinetic value $\tau_{\rm k}$, the slip distance must be of the order of the size $D$ of the macroasperity contact regions. If, due to elasticity, different regions at the interface have already slipped a distance larger than $\lambda_{\rm c}\approx D$ when full slip occurs, then the breakloose friction will be smaller than $\tau_{\rm s}A$, which would be the friction force if all contact regions slipped simultaneously, as in the case of a rigid solid.
For the situation in Fig. 27, one can show that the elastic deformations are characterized by a length scale $\lambda_{\rm el}\approx(\tau_{\rm s}-\tau_{\rm k})L^{2}/Eh$, where $E$ is the Young’s modulus, $h$ is the block height, and $L$ is its length in the sliding direction. If $\lambda_{\rm el}>\lambda_{\rm c}$, the breakloose friction force is smaller than $\tau_{\rm s}A$, and if $\lambda_{\rm el}\gg\lambda_{\rm c}$, it may vanish.
@@ 7 Triangular slider on rubber surfaces
$p(x)=p_{0}\,{\rm cosh}^{-1}\left(\frac{a}{x}\right)$ (34)
@@
$\tau_{\rm f}(v)=[\mu(v)-\mu_{\rm visc}(v)]{F_{\rm N}\over w(v)L_{y}}$ (35)
Here, $w(v)$ is the width of the contact region, $\mu(v)$ is the total friction coefficient, and $\mu_{\rm visc}$ is the viscoelastic contribution, as given by the green and blue curves in Fig. 42. Using this equation gives a frictional shear stress $\tau_{\rm f}(v)$ that increases nearly linearly with the logarithm of the sliding speed, as expected for thermally activated processes in the low-velocity regime.
@@ 8 Lubricated rubber friction
This review mainly focuses on rubber friction on dry surfaces. However, rubber friction on wet surfaces, or on surfaces covered by lubrication or contamination films, is equally important. Applications include tires on wet or contaminated road surfaces and dynamic seals with surfaces covered by oil films. Other examples are the friction between shoes and contaminated ground surfaces, and syringes where the interface between the rubber stopper and the glass or polymer barrel is often lubricated with silicone oil. Here we will not present a complete study of rubber friction on lubricated surfaces, but we highlight a few basic facts.
When an elastic body slides on a substrate in the presence of a fluid, one can distinguish three regimes as indicated in Fig. 45, which schematically shows the friction coefficient as a function of the sliding speed. Here we focus on the behavior of thin fluid films that may exist in the boundary and early mixed lubrication regions of the Stribeck curve. Thin films between solid surfaces may be stable or unstable depending on interfacial energies and the nominal contact pressure. We first consider smooth surfaces as studied in Refs. [68, 69, 70, 71, 72].
@@
$v_{\rm c}\sim{|S|\over\eta}\left({|S|\over ER}\right)^{1/3},$
@@
The surface tension of water (and glycerol) is approximately $0.07\ {\rm J/m^{2}}$. If $S$ is positive, its magnitude is expected to be at most a few times $0.01\ {\rm J/m^{2}}$. Using $S=0.01\ {\rm J/m^{2}}$ and $d^{*}=1\ {\rm nm}$ gives $p^{*}=10\ {\rm MPa}$, which is also the typical pressure in the rubber-concrete real contact areas. Assuming that the relative asperity contact area $A/A_{0}$ is approximately $0.01$ (as estimated in Ref. [75]), and taking $p_{0}=0.12\ {\rm MPa}$, we find that $p=p_{0}A_{0}/A=12\ {\rm MPa}$. Hence, the fluid film may be removed from the asperity contact regions even if $S>0$, provided that $S$ is small enough.
For sliding speeds $v>0.1\ {\rm mm/s}$, the friction in the fluids is smaller than in the dry state. However, a very thin fluid film is sufficient to explain the observed drop in the friction coefficient. For example, if the fluid film is approximately $3\ {\rm nm}$ thick, then for pure glycerol ($\eta\approx 1.4\ {\rm Pa\cdot s}$, not accounting for shear thinning) at the sliding speed $v=1\ {\rm cm/s}$, the shear stress is expected to be $\tau_{\rm f}\approx\eta v/d\approx 5\ {\rm MPa}$. This gives a friction coefficient of approximately $\tau_{\rm f}A/p_{0}A_{0}\approx 0.2$, which is consistent with experimental observations.
We therefore propose that the reduction in the friction is due to forced wetting, where the effective spreading pressure $S^{*}$ is positive only as a result of the external pressure. As discussed above, if $S$ is positive, a thin fluid film will separate the surfaces of the solids in the absence of external forces. This remains true even under an applied squeezing pressure $p_{0}$, provided it is not too high. However, if $S$ is small or the squeezing pressure is high enough, the fluid film may be removed by forced interfacial dewetting.

### v3

@@ 5.1 Analytical theory of rubber friction on rough surfaces
For the relatively stiff rubber compounds used in practical applications such as tires and seals, there is consistent experimental evidence that crack-opening mechanisms contribute little to the overall sliding or rolling friction. This is primarily due to the multiscale roughness of the surface, which strongly suppresses adhesion [61]. In Sec. 8it is shown that for rubber compounds which exhibiting strong adhesion on smooth surfaces under dry conditions, but no adhesion in water, display nearly identical sliding friction on rough surfaces under both dry and wet conditions [26]. Similar observations have been made by Greenwood and Tabor [18].
Figure 21: The friction coefficient, divided by the surface rms slope $\xi$, as a function of the dimensionless sliding speed $q_{0}v\tau$, for a surface with the height profile $h_{0}\sin(q_{0}x)$ (dashed lines), and for a surface with one-dimensional self-affine roughness, with the ratio between the large and small wavenumber cut-offs $q_{1}/q_{0}=60$ and the Hurst exponent $H=0.8$ (solid lines). The viscoelastic modulus is of the form given in (4), with $E_{1}/E_{0}=10$ and relaxation time $\tau$. The applied pressure is $p=0.075\,p_{0}$ in (a) and $0.18\,p_{0}$ in (b), where the reference pressure is $p_{0}=E_{0}^{*}q_{0}h_{\rm max}/2$, and $h_{\rm max}$ is the height of the highest asperity relative to the average surface plane. Adapted from Ref. [62].
The influence of adhesion on the sliding friction for a viscoelastic solid with the simple modulus given by (4), with $E_{0}=10\ {\rm MPa}$ and $E_{1}/E_{0}=10$, was studied theoretically in Ref. [62]. The sliding friction was calculated for a surface with the height profile $h_{0}\sin(q_{0}x)$, and for a surface with one-dimensional (1D) self-affine roughness, with the ratio between the large and small wavenumber cut-offs $q_{1}/q_{0}=60$ and the Hurst exponent $H=0.8$. The interfacial work of adhesion was $\gamma=0.01\ {\rm J/m^{2}}$, as is typical for rubber in contact with many solids (smooth surfaces).
Fig. 21 shows the friction coefficient $\mu$ divided by the surface rms slope $\xi$ as a function of the dimensionless sliding velocity $v/v_{0}$ for two different values of the dimensionless pressure $p/p_{0}$. The reference velocity is $v_{0}=1/(q_{0}\tau)$, and the reference pressure is $p_{0}=E_{0}^{*}q_{0}h_{\rm max}/2$, where $h_{\rm max}$ is the height of the highest asperity relative to the average surface plane.
Results are shown both with and without adhesion. As the pressure increases, the influence of adhesion decreases. It is not clear how these results would be modified for real surfaces with 2D roughness and with larger values of $q_{1}/q_{0}$.
@@ 5.2 Comparing the theory with experimental results
The results in Fig. 23 demonstrate that, in most cases, at low sliding speeds, water is almost completely expelled from the asperity contact regions, resulting in a friction coefficient similar to that observed under dry conditions. However, at much higher sliding speeds, there is insufficient time for the water film to be fully squeezed out, leading to a significant reduction in friction. At very high speeds, hydroplaning may occur, and the friction becomes negligible.
@@
One important quantity that characterizes rubber is its glass transition temperature, $T_{\rm g}$, defined as the temperature at which ${\rm tan}\,\delta$ reaches its maximum for a given reference frequency (we use $\omega=0.01\ {\rm s}^{-1}$). Due to the WLF temperature-frequency relationship, $T_{\rm g}$ also provides information about the frequency at which ${\rm tan}\,\delta$ is maximal.
@@
Adhesion and viscoelasticity can have a strong influence on the area of real contact [65, 66]. The stiffening of the elastic properties of rubber when perturbed at high frequencies results in a drastic decrease in the contact area with increasing sliding speed in a region of intermediate velocities [see Fig. 18(b)]. At very high sliding speeds, the contact area can increase due to frictional heating, which reduces the elastic modulus of the rubber.
At low sliding speeds on smooth surfaces, experimental observations also indicate that adhesion can increase the contact area with increasing sliding speed [65] [see Fig. 27]. This may be due to the increased energy required to propagate opening cracks at higher sliding speeds.
Figure 27: Two contact footprints between a Buna-N nitrile rubber sphere (diameter $4.8\ {\rm mm}$, with rms roughness $h_{\rm rms}=5.2\ {\rm\mu m}$) and a borosilicate float glass optical window. (a) shows the contact area before sliding, and (b) at a sliding speed of $20\ {\rm\mu m/s}$. The contact area during sliding is larger by a factor of $49/39\approx 1.26$.
Figure 28: Master curve of the friction coefficient for acrylonitrile-butadiene rubber on smooth but macroscopically wavy glass. The reference temperature is $T_{0}=20^{\circ}{\rm C}$. The master curve was obtained by shifting velocity segments (velocity range approximately from $1\ {\rm\mu m/s}$ to $1\ {\rm cm/s}$), measured at temperatures between $15^{\circ}$ and $85^{\circ}{\rm C}$, using the bulk viscoelastic shift factor $a_{T}$. When the friction coefficient decreases with increasing sliding speed, stick-slip occurs. For these velocities, the reported friction coefficient corresponds to the peak friction force during the stick-slip oscillations. Adapted from Ref. [19].
Figure 29: (a) Shear stress as a function of velocity and temperature for sliding between PDMS ($M=2.7\ {\rm kg/mol}$) and a silicon wafer coated with a hexadecylsiloxane self-assembled monolayer (SAM). Open, gray, and black circles represent data measured at 298, 318, and 348 K, respectively. (b) The temperature-dependent shear stress data shifted to room temperature using an Arrhenius shift factor. The activation energy for sliding between PDMS and the SAM is estimated to be $25\ {\rm kJ/mol}$. Open, gray, and black circles again represent data measured at 298, 318, and 348 K, respectively.
Sliding friction on smooth surfaces
Rubber friction for tire tread compounds on smooth surfaces is a complicated topic, because mobile components (waxes), which are added to the rubber compounds, diffuse to the rubber surface and are slowly sheared off during sliding, resulting in a friction coefficient that depends sensitively on the sliding distance. Experiments on “simpler” rubber compounds such as poly(dimethylsiloxane) (PDMS) elastomers give more reproducible results.
Grosch [19] studied the friction of acrylonitrile-butadiene rubber on smooth but macroscopically wavy glass. Fig. 28 shows the master curve of the friction coefficient at the reference temperature $T_{0}=20^{\circ}{\rm C}$. The master curve was obtained by shifting velocity segments measured at temperatures between $15^{\circ}$ and $85^{\circ}{\rm C}$, using the bulk viscoelastic shift factor $a_{T}$. The friction master curve has the general form expected from the Schallamach model [16] of the adhesive contribution to friction, but the width of the curve is larger than predicted by the Schallamach model and instead agrees better with the prediction of the Persson-Volokitin model [21].
Vorvolakos and Chaudhury [67] performed experiments in which spherical cups of PDMS were slid on Si wafers coated with a hexadecylsiloxane monolayer, or on thin films of polystyrene. Elastomers with various molecular weights were used on both surfaces. Fig. 29 shows the shear stress as a function of velocity and temperature between PDMS and a Si wafer with a self-assembled monolayer of hexadecylsiloxane. From the temperature-dependent studies, the activation energy of friction on both surfaces was estimated to be $25\ {\rm kJ/mol}$, which compares well with values obtained from studies of melt dynamics.
The friction first increases with sliding velocity, reaches a maximum, and then decreases, as expected from the Schallamach model. The maximum of the frictional shear stress $\sigma^{*}$ was found to be roughly proportional to the square root of the elastic modulus, as expected from the Schallamach model if the size of the stress domains is independent of the modulus [see ()]. The sliding velocity $v^{*}$ at which $\sigma_{\rm f}$ is maximum differs for the two substrates, which may reflect different relaxation times $\tau_{0}$, as expected since the free energy barrier for attachment depends on the substrate.
Figure 30: The frictional shear stress divided by the elastic modulus raised to the $1/2$ power as a function of the normalized sliding speed $v/v^{*}$ (log-log scale). The data points are for PDMS rubber with different moduli sliding on three different surfaces (see inset). Adapted from Ref. [68].
Maksuta et al. [68] measured the frictional shear stress for PDMS rubbers with different elastic moduli sliding on three surfaces with similar interfacial chemistry but different roughness. Analysis of the experiments showed that friction on all the surfaces was dominated by adhesive contributions rather than viscoelastic dissipation, and that in all cases the rubber made complete contact with the substrate.
Contrary to the expectation that roughness only influences the available contact area for adhesive friction, it was found that roughness caused a shift in the velocity at which the adhesive friction is maximum. The authors suggested that this resulted from roughness-induced vertical oscillations of the rubber surface during sliding. The frictional shear stress scaled with the elastic modulus as $E^{1/2}$ (see Fig. 30), as expected from the Schallamach model for adhesive friction on smooth surfaces.
@@ 5.3 Dynamical friction
Figure 34: Slip at the onset of sliding when the driving force is applied on one side of the rubber block. The region where slip occurs gradually expands as the system approaches the onset of global slip, at which point the entire contact area is sliding. The frictional shear stress in regions where the slip displacement $u>\lambda_{\rm asp}$ is equal to the kinetic frictional shear stress $\sigma_{\rm k}$. Consequently, the breakloose friction force is smaller than $F_{\rm s}=A\sigma_{\rm s}$ and approaches $F_{\rm k}=A\sigma_{\rm k}$ as the elastic modulus decreases.
@@
Consider a “kinetic” frictional shear stress $\sigma_{\rm k}$ acting on the contact area $A$ between a block and substrate during sliding. After a stationary contact duration $t_{\rm s}$, the shear stress at the new onset of sliding becomes $\sigma_{\rm s}$, which we assume to be larger than $\sigma_{\rm k}$, e.g., due to thermally activated interfacial bond formation. To reduce the shear stress to its kinetic value $\sigma_{\rm k}$, the slip distance must be of the order of the size $D$ of the macroasperity contact regions. If, due to elasticity, different regions at the interface have already slipped a distance larger than $\lambda_{\rm c}\approx D$ when full slip occurs, then the breakloose friction will be smaller than $\sigma_{\rm s}A$, which would be the friction force if all contact regions slipped simultaneously, as in the case of a rigid solid.
For the situation in Fig. 34, one can show that the elastic deformations are characterized by a length scale $\lambda_{\rm el}\approx(\sigma_{\rm s}-\sigma_{\rm k})L^{2}/Eh$, where $E$ is the Young’s modulus, $h$ is the block height, and $L$ is its length in the sliding direction. If $\lambda_{\rm el}>\lambda_{\rm c}$, the breakloose friction force is smaller than $\sigma_{\rm s}A$, and if $\lambda_{\rm el}\gg\lambda_{\rm c}$, it may vanish.
@@
Many other experimental and theoretical studies have shown that the breakloose (or static) friction force depends on the geometry of the contacting bodies and on how the driving force is applied. As an example, Zhang and Ciavarella [70] (see also Refs. [71, 72]) used the Cattaneo-Mindlin friction model with Griffith friction (where the effect of roughness is included in a fracture energy linearly dependent on the local pressure), and predicted that at low normal loads, the breakloose friction coefficient depends on the geometry of the contact.
@@ 6 Rolling friction for a hard cylinder and sphere on a viscoelastic solid
Adhesive contribution to rolling friction
The theory presented above for rolling friction assumes negligible adhesion. However, for very smooth and clean surfaces adhesion is important, and for small normal forces it can give the dominant contribution to the rolling friction force.
Consider a rigid sphere or cylinder in contact with a viscoelastic solid with a flat surface. When adhesion occurs, the boundary line between the contact and non-contact surface area can be considered as an adhesive crack tip. When a cylinder (or sphere) rolls on a rubber surface, the rubber comes into contact with the cylinder on one side (closing crack) and moves out of contact on the other side (opening crack). If a crack moves with a finite speed $v$, the energy required to break the bonds at the crack tip is in general larger at the opening crack than the energy gained by bond formation at the closing crack. This adhesion hysteresis can be understood by a simple example.
Figure 46: (a) A solid with a flat surface (black) approaching a solid consisting of long-chain molecules. The solids make contact only when they are at atomic separation. (b) During separation, a chain molecule is bonded to the upper solid, and the bond breaks only when the chain is nearly fully stretched.
Consider a flat surface approaching a solid consisting of long-chain molecules, see Fig. 46. At a finite approach speed $v$, the solids make contact only when they are within the range of the interaction potential between the solids (usually of order 1 nm or less). This bond formation gives an energy gain per unit surface area denoted by $\gamma_{0}$. Next, we separate the solids. If the separation velocity is high, the bond between the chain molecules and the flat surface breaks only when the chain molecules are nearly fully stretched. Hence, the work per unit surface area $\gamma_{1}$ required to separate the surfaces may be much larger than $\gamma_{0}$. In general, $\gamma_{0}(v)$ and $\gamma_{1}(v)$ depend on the speed $v$. For infinitesimally slow motion ($v\rightarrow 0$), thermal equilibrium occurs at all times and $\gamma_{0}(0)=\gamma_{1}(0)$.
Figure 47: (a) A hard cylinder or ball rolling on a flat rubber track. The regions where most of the viscoelastic energy dissipation occurs are indicated by the red dashed areas. (b) A rigid cylinder rolling down a rubber track. From the tilt angle $\theta$ one can calculate the effective interfacial energy $\gamma_{\rm eff}$.
For viscoelastic solids, the opening crack propagation energy $\gamma_{\rm open}(v)$ usually increases rapidly with increasing crack-tip velocity due to viscoelastic energy dissipation close to the crack tip [see Fig. 47(a)] [77]. Hence, $\gamma_{\rm open}(v)$ can be much larger than $\gamma_{1}(v)$. Similarly, viscoelastic energy dissipation close to the closing crack tip reduces the energy gain from bond formation, so that $\gamma_{\rm close}(v)$ may be much smaller than $\gamma_{0}(v)$. Rolling friction depends on the difference $\gamma_{\rm open}(v)-\gamma_{\rm close}(v)\approx\gamma_{\rm open}(v)$ (see below), so no direct information about $\gamma_{\rm close}(v)$ can be obtained from rolling friction experiments. However, both $\gamma_{\rm open}(v)$ and $\gamma_{\rm close}(v)$ can be deduced from adhesion experiments in which a spherical ball is moved into and out of contact with the viscoelastic solid. One example is shown in Fig. 48, where negligible adhesion occurs on approach (closing crack), while strong adhesion is observed on retraction.
Figure 48: A glass sphere (radius $R=2\ {\rm cm}$) brought into contact with a weakly crosslinked PDMS rubber surface, kept in contact for about 1 hour, and then removed. The approach and retraction speed is $0.87\ {\rm\mu m/s}$. Note that no adhesion force can be observed during contact formation (closing crack) (dashed rectangle), while a large attractive force is observed during retraction (opening crack). The pull-off force is related to $\gamma_{\rm open}$ via the JKR formula $F_{\rm pull-off}=(3\pi/2)\gamma_{\rm open}$. Adapted from Ref. [78].
The adhesive contribution to the rolling friction of a cylinder on a flat viscoelastic solid can be obtained as follows: if the cylinder has length $L$ and rolls a distance $\Delta x$, the opening and closing crack tips move over the surface area $A=L\Delta x$, so the energy dissipated at the opening crack tip is $\gamma_{\rm open}(v)L\Delta x$ and the energy gained at the closing crack tip is $\gamma_{\rm close}(v)L\Delta x$, giving the net energy loss
$\Delta E=[\gamma_{\rm open}(v)-\gamma_{\rm close}(v)]L\Delta x$
If $f_{\rm R}$ is the rolling force per unit length, the rolling friction energy is $f_{\rm R}L\Delta x$, giving
$f_{\rm R}=\gamma_{\rm open}(v)-\gamma_{\rm close}(v)=\gamma_{\rm eff}(v)$
Note that $f_{\rm R}$ is independent of the normal force, so the rolling friction coefficient $\mu=f_{\rm R}/f_{\rm N}$ increases without limit as $f_{\rm N}\rightarrow 0$. This differs from the viscoelastic contribution studied before [see (30)], where $f_{\rm R}$ depends nonlinearly on $f_{\rm N}$ (typically $f_{\rm R}\sim f_{\rm N}^{1/2}$ for a cylinder and $F_{\rm R}\sim F_{\rm N}^{1/3}$ for a sphere), and in particular $f_{\rm R}\rightarrow 0$ as $f_{\rm N}\rightarrow 0$. Hence, the adhesive contribution dominates the rolling friction for sufficiently small normal loads. In Ref. [79], an equation was derived for the load where the rolling friction crosses over from the adhesive to the viscous limiting case:
$\gamma_{\rm eff}\approx\left({\sigma_{0}\over E_{0}}\right)^{3}R{\rm Im}E(\omega_{0})$ (34)
where $\sigma_{0}$ is the average contact pressure, and $\omega=r/d$, where $d$ is the width of the contact regions. $\sigma_{0}$ and $d$ can be estimated using Hertz contact theory. (34) shows that for large enough contact pressure the viscoelastic bulk contribution will dominate.
Figure 49: The logarithm (base 10) of the effective interfacial energy $\gamma_{\rm eff}(v)$ as a function of the logarithm of the crack-tip velocity $v$, deduced from rolling experiments with glass cylinders on smooth plane rubber tracks. The different symbols correspond to different glass cylinder loads (0, 66, $18$, $28$, and $102\ {\rm g}$), where increasing velocity corresponds on average to increasing load. Results are shown for natural rubber and polychloroprene rubber tracks. Adapted from Ref. [80].
The rolling friction can be measured using a tilted substrate, as indicated in Fig. 47(b). The tangential component of the gravitational force acting on the cylinder is $F=mg\sin\theta$, which must equal the friction force $Lf_{\rm R}$ during rolling at constant speed. This gives $\gamma_{\rm eff}(v)=(mg/L)\sin\theta$. We note that steady rolling occurs only if $\gamma_{\rm eff}(v)$ increases with the crack-tip velocity $v$, which is almost always the case for rubber due to the strong increase in viscoelastic energy dissipation at the crack tip with increasing velocity.
Roberts and Thomas [80] performed rolling friction experiments with glass cylinders on smooth plane rubber tracks. Fig. 49 shows $\gamma_{\rm eff}(v)$ for natural rubber and polychloroprene as a function of the crack-tip velocity $v$ (log-log scale). The different symbols correspond to different glass cylinder loads (0, 66, $18$, $28$, and $102\ {\rm g}$). As the cylinder mass increases, the rolling velocity increases, and the fact that all data points collapse onto a common master curve shows that the rolling friction is independent of the normal force and depends only on the rolling speed.
@@ 7 Triangular slider on rubber surfaces
$p(x)=p_{0}\,{\rm cosh}^{-1}\left(\frac{a}{x}\right)$
@@
$\sigma_{\rm f}(v)=[\mu(v)-\mu_{\rm visc}(v)]{F_{\rm N}\over w(v)L_{y}}$ (35)
Here, $w(v)$ is the width of the contact region, $\mu(v)$ is the total friction coefficient, and $\mu_{\rm visc}$ is the viscoelastic contribution, as given by the green and blue curves in Fig. 53. Using this equation gives a frictional shear stress $\sigma_{\rm f}(v)$ that increases nearly linearly with the logarithm of the sliding speed, as expected for thermally activated processes in the low-velocity regime.
@@
In Ref. [83], the authors studied the sliding of non-cylindrical wedges on a viscoelastic half-space. Using the boundary element method, they found that for wedges with the height profile $h(x)\sim|x|^{k}$, with $k=4,6,$ and $8$, the viscoelastic contribution to the friction force differs from the prediction of (30). However, if the friction coefficients are scaled by $E_{0}/p_{0}$, where $E_{0}$ and $p_{0}$ are the modulus and mean pressure at zero speed, nearly identical results are obtained for all wedges, including the cylindrical wedge (where $h(x)\sim x^{2}$). Hence, the friction coefficients as a function of sliding speed can be obtained from the cylindrical case, where (30) is accurate, by scaling.
@@ 8 Lubricated rubber friction
This review mainly focuses on rubber friction on dry surfaces. However, rubber friction on wet surfaces, or on surfaces covered by lubrication or contamination films, is equally important. Applications include tires operating on wet or contaminated road surfaces and dynamic seals in which the sliding surfaces are covered by oil films. Other examples include the friction between shoes and contaminated ground surfaces, and syringes, where the interface between the rubber stopper and the glass or polymer barrel is often lubricated with silicone oil.
Here we do not present a comprehensive study of rubber friction on lubricated surfaces, but we highlight a few basic facts.
When an elastic body slides on a substrate in the presence of a fluid, one can distinguish three regimes, as illustrated in Fig. 56, which schematically shows the friction coefficient as a function of sliding speed. Here, we focus on the behavior of thin fluid films that may exist in the boundary and early mixed lubrication regions of the Stribeck curve.
The hydrodynamic region prevails when there is no direct contact between the solids. For randomly rough surfaces with a roll-off, the highest asperities are typically on the order of a few times $h_{\rm rms}$. However, if the surfaces have already been in contact in the boundary lubrication region, the highest asperities may be significantly reduced due to deformation or wear.
In the boundary lubrication region, most of the load is carried by the area of real contact. Nevertheless, a small fraction of the load—typically around 10%—may be supported by pressurized fluid trapped in sealed-off regions.
Thin films between solid surfaces may be stable or unstable depending on interfacial energies and the nominal contact pressure. We first consider smooth surfaces as studied in Refs. [84, 85, 86, 87, 88].
@@
$v_{\rm c}\sim{|S|\over\eta}\left({|S|\over ER}\right)^{1/3},$ (38)
@@
The surface tension of water (and glycerol) is approximately $0.07\ {\rm J/m^{2}}$. If $S$ is positive, its magnitude is expected to be at most a few times $0.01\ {\rm J/m^{2}}$. Using $S=0.01\ {\rm J/m^{2}}$ and $d^{*}=1\ {\rm nm}$ gives $p^{*}=10\ {\rm MPa}$, which is also the typical pressure in the rubber-concrete real contact areas. Thus, assuming that the relative asperity contact area $A/A_{0}$ is approximately $0.01$ (as estimated in Ref. [91]), and taking $p_{0}=0.12\ {\rm MPa}$, we find that $p=p_{0}A_{0}/A=12\ {\rm MPa}$. Hence, the fluid film may be removed from the asperity contact regions even if $S>0$, provided that $S$ is small enough.
For sliding speeds $v>0.1\ {\rm mm/s}$, the friction in the fluids is smaller than in the dry state. However, a very thin fluid film is sufficient to explain the observed drop in the friction coefficient. For example, if the fluid film is approximately $3\ {\rm nm}$ thick, then for pure glycerol ($\eta\approx 1.4\ {\rm Pa\cdot s}$, not accounting for shear thinning) at the sliding speed $v=1\ {\rm cm/s}$, the shear stress is expected to be $\sigma_{\rm f}\approx\eta v/d\approx 5\ {\rm MPa}$. This gives a friction coefficient of approximately $\sigma_{\rm f}A/p_{0}A_{0}\approx 0.2$, which is consistent with experimental observations.
The drop in the friction coefficient in glycerol, starting at the sliding speed $v\approx 0.03\ {\rm mm/s}$, is likely due to forced wetting. Assuming $|S^{*}|\approx 0.01\ {\rm J/m^{2}}$, a fluid viscosity of $\eta\approx 1\ {\rm Pas}$, an elastic modulus $E=10^{7}\ {\rm Pa}$, and using a macroasperity radius of curvature $R=26\ {\rm mm}$—as obtained from the concrete surface roughness power spectrum (see Ref. [91])—we obtain from (38) a critical velocity $v_{\rm c}=0.03\ {\rm m/s}$, which is consistent with the results shown in Fig. 60.